# MRAMG-Bench: A BeyondText Benchmark for Multimodal Retrieval-Augmented Multimodal Generation

### v1

@@ Abstract.
Recent advancements in Retrieval-Augmented Generation (RAG) have shown remarkable performance in enhancing response accuracy and relevance by integrating external knowledge into generative models. However, existing RAG methods primarily focus on providing text-only answers, even in multimodal retrieval-augmented generation scenarios. In this work, we introduce the Multimodal Retrieval-Augmented Multimodal Generation (MRAMG) task, which aims to generate answers that combine both text and images, fully leveraging the multimodal data within a corpus. Despite the importance of this task, there is a notable absence of a comprehensive benchmark to effectively evaluate MRAMG performance. To bridge this gap, we introduce the MRAMG-Bench, a carefully curated, human-annotated dataset comprising 4,346 documents, 14,190 images, and 4,800 QA pairs, sourced from three categories: Web Data, Academic Papers, and Lifestyle. The dataset incorporates diverse difficulty levels and complex multi-image scenarios, providing a robust foundation for evaluating multimodal generation tasks. To facilitate rigorous evaluation, our MRAMG-Bench incorporates a comprehensive suite of both statistical and LLM-based metrics, enabling a thorough analysis of the performance of popular generative models in the MRAMG task. Besides, we propose an efficient multimodal answer generation framework that leverages both LLMs and MLLMs to generate multimodal responses. Our datasets are available at: https://huggingface.co/MRAMG.
Multimodal Retrieval-Augmented Multimodal Generation, Large Language Model, Multimodal Large Language Model, Evaluation
journalyear: 2024copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Information systems Information retrieval
@@ 1. Introduction
With the evolution of generative models, the shift from traditional LLMs to multimodal large language models (MLLMs) has facilitated a more comprehensive understanding by integrating both textual and visual inputs (Wang et al., 2024b, a). Correspondingly, Multimodal Retrieval-Augmented Generation (MRAG) (Chen et al., 2022) builds upon traditional RAG by incorporating both textual and visual information, thereby enhancing the quality of generated responses through the image understanding capabilities of MLLMs. However, while the retrieval process incorporates multimodal information, current MRAG methods primarily focus on multimodal input and text-based output. Despite their advanced image comprehension capabilities, MLLMs frequently exhibit hallucinations when converting visual inputs into textual outputs (Bai et al., 2024). This raises a natural question:“Why not present the images directly rather than describing their content?
@@
  • Images as a catalyst for richer answer: “A picture is the icing on the cake.” In applications such as tourist spot descriptions, integrating both textual and visual content can significantly enhance the quality of the generated response.
@@
To address this critical demand, we introduce the Multimodal Retrieval-Augmented Multimodal Generation (MRAMG) task, which aims to fully leverage the rich multimodal information within a corpus by generating integrated text-and-image answers (See Section 3 for the definition). However, current benchmarks for evaluating this critical task suffer from a notable lack of suitable datasets and scientifically rigorous evaluation metrics, as shown in Table 1. This significant shortcoming hinders progress in this emerging field.
@@
To bridge this gap, we introduce the MRAMG-Bench, a novel benchmark designed to evaluate the MRAMG task comprehensively. The MRAMG-Bench consists of six carefully curated datasets, comprising 4,346 documents, 14,190 images, and 4,800 QA pairs, sourced from three domains—Web Data, Academic Paper Data, and Lifestyle Data—across seven distinct data sources. Furthermore, MRAMG-Bench introduces unique challenges, such as hierarchical difficulty levels and order-based reasoning, which rigorously test the reasoning capabilities of both LLMs and MLLMs. To the best of our knowledge, MRAMG-Bench is the first benchmark that challenges models to autonomously determine the number, selection, and ordering of images in their responses, effectively simulating the complex scenarios encountered in real-world user interactions.
@@
This benchmark is systematically constructed through a multi-stage workflow with both automatic construction of GPT-4o and meticulous formulation of human annotators, which is illustrated in Figure 2. Traditional evaluations of multimodal answers often rely on subjective metrics, such as image-text consistency, which are assessed through LLM-based judgments or human evaluations (Zhang et al., 2023; Zhu et al., 2024a). However, these methods suffer from high subjectivity or evaluator inconsistency. To mitigate these issues, we introduce a statistically grounded evaluation framework that systematically assesses both retrieval and generation performance. By integrating objective statistical metrics with LLM-based assessments, our framework delivers a comprehensive, multi-dimensional evaluation of multimodal answers, ensuring a fair and rigorous assessment process.
Additionally, we propose a general multimodal answer generation framework that integrates rule-based and model-based approaches. This flexible and scalable solution enables both LLMs and MLLMs to generate interleaved text-image responses. Our main contributions can be summarized as follows:
@@
  • We formalize the MRAMG task , a benchmark consisting of 4,800 human-annotated instances across multiple domains, featuring hierarchical difficulty levels and order-based reasoning challenges. To ensure rigorous evaluation, we propose a robust, statistically grounded set of metrics that assess multimodal responses across various performance dimensions.
@@
  • We propose a flexible multimodal answer generation framework that leverages LLMs and MLLMs, combining rule-based and model-based approaches.
@@ 3. Task Formulation
In this section, we formally introduce the definition of the task in Multimodal Retrieval-Augmented Multimodal Generation (MRAMG), focusing on constructing multimodal answers $\mathcal{A}$ based on a given text-form user query $q$ and its associated multimodal information $\mathcal{D}$. See Figure 1 for an illustration.
@@ 4. Dataset Construction
Figure 2. The construction pipeline of MRAMG-Bench. It is divided into three stages: (1) Data Selection and Preprocessing, we collect, clean and preprocess data to ensure the quality of initial datasets. (2) QA Generation and Refinement, we formulate questions and answers and refine the formulated QA pairs. (3) Data Quality Check, we engage annotators and experts in a three-stage data quality check to ensure high-quality of the entire benchmark.
In this section, we detail the construction of MRAMG-Bench. It consists of three domains: Web, Academic Paper and Lifestyle, which respectively contains web-based datasets—MRAMG-Wit, MRAMG-Wiki, MRAMG-Web, and a scientific dataset-MRAMG-Arxiv, and two lifestyle-oriented datasets-MRAMG-Recipe, MRAMG-Manual. The process of creating datasets in different domains can be divided into three stages: (1) Data Selection and Preprocessing, (2) QA Generation and Refinement, (3) Data Quality Check. The overview of this process is shown in Figure 2.
@@ 4.1. Data Selection and Preprocessing
We undertake three steps during this stage: (1) Collect, (2) Filter Data and Enhance Context, and (3) Refine.
4.1.1. Collect
@@
This category includes Wikipedia pages and articles, derived from Wit dataset (Srinivasan et al., 2021), WikiWeb2M dataset (Burns et al., 2023), and WebQA dataset (Chang and Bisk, 2022). These datasets typically feature a rich integration of text and images, with the insertion order of images in Wikipedia, which serves as a reliable reference. We select val sets and test sets from Wit dataset (Srinivasan et al., 2021), and sample 82k entries from the 100k entries in the test set of WikiWeb2M dataset (Burns et al., 2023) , and choose img-posFacts set in WebQA dataset (Chang and Bisk, 2022), as the initial collection of MRAMG-Wit dataset, MRAMG-Wiki dataset and MRAMG-Web dataset. These datasets primarily consist of single-image documents with relatively short length. Consequently, we categorize the three Web Data datasets as level 1: easy (difficulty level), where models can easily handle our task in this domain due to simple scenarios.
Academic Paper Data.
@@
Datasets in this domain features in fine-integration of a sequence of operation images and instruction texts. These dataset are thoroughly applicable to practical scenarios encountered in users’ daily lives, which are collected from RecipeQA (Yagcioglu et al., 2018), ManualsLib [^1] https://www.manualslib.com and the Technical Dataset [^2] https://www.kaggle.com/datasets/ahmedhgabr/technical-illustration/data. We select the evaluation set from RecipeQA (Yagcioglu et al., 2018), and select 40 categories of instruction manuals with rich text-image content for subsequent construction of MRAMG-Recipe dataset and MRAMG-Manual dataset, respectively. We define Lifestyle Data as level 3: difficult, as they exhibit a significantly higher density of images and encompass diverse problem types, including single-image, no-image, and multi-image answers. Moreover, these datasets pose a significant challenge for models in determining the order of image insertion.
4.1.2. Filter Data and Enhance Context
@@
For Academic Paper Data, we engage ten graduate students specializing in computer science, with expertise in fields such as Natural Language Processing (NLP) and Computer Vision (CV), to serve as expert annotators. Each annotator is assigned papers within their specific domain of expertise, ensuring precise and contextually relevant annotations. The annotators meticulously read the selected papers, identify and extract textual content and corresponding figures that are most suitable for creating interleaved text-image contexts. This ensures accurate alignment of text and images in contexts.
For Lifestyle Data, in MRAMG-Recipe dataset, each recipe contains multiple step texts and each step text contains multiple images. We first combine multiple images into one image for each recipe step text and insert them at the end of each step text. We then aggregate all the recipe step texts as an entire image-text interleaved recipe context. In MRAMG-Manual dataset, we employ MinerU (Wang et al., 2024c) to parse the PDF manual documents into markdown format, preserving the original contextual order of images and text within each manual. A team of annotators is tasked with removing sensitive or irrelevant information, adjusting image placements within the content, and correcting any errors in the text.
4.1.3. Refine
@@ 4.2. QA Generation and Refinement
This stage can be further divided into three steps: (1) question generation, (2) answer generation, (3) QA pair refine.
@@ 4.2.1. Question Generation
For MRAMG-Web dataset, we retain the questions contained in original dataset. As for MRAMG-Wiki, MRAMG-Wit and MRAMG-Recipe datasets, we leverage GPT-4o to generate high-quality, context-specific questions tailored to the provided text and images. Moreover, the questions in MRAMG-Manual are generated manually by human annotators. As for MRAMG-Arxiv dataset, we engage expert annotators to meticulously formulate questions. Overall, We adhere to follow two criteria: (1) The questions are constructed based on the corresponding contexts. (2) The questions must be natural and practical and they present potential to be answered with texts integrated with images.
@@ 4.2.2. Answer Generation
For Web Data, given the generated or original questions $Q$ and corresponding context $C$, we generate image-text interleaved answers $A$ using GPT-4o following CoT reasoning strategy (Wei et al., 2022): (1) Question Validity Assessment, we first ensure valid questions $ValQ$ can be directly answered by contexts $C$ and remove invalid questions $InvalQ$. (2) Evidence Extraction, we then extract evidences $E$ from contexts $C$ to substantiate the answer $A$ for later answer construction. (3) Answer Construction, we construct highly reliable, accurate and coherent answer based on valid questions $ValQ$ and evidences $E$. (4) Image Integration we finally enrich the constructed answers $A$ with a series of images integrated in the optimal positions.
For Academic Paper Data, the expert annotators are responsible for reviewing previous formulated questions, constructing answers where images or tables play an integral role, and seamlessly integrating visual elements into the answer, to provide a cohesive and multimodal understanding of the content. For Lifestyle Data, in MRAMG-Recipe dataset, there are two types of questions: questions related to the entire recipe (recipe-specific) and questions only relevant to a specific step (step-specific). We combine all the step texts as the answers of recipe-specific questions and use answers generated by GPT-4o as the answers related to step-specific questions. We manually insert images at the end of answer texts. In MRAMG-Manual dataset, answers are meticulously formulated by annotators, to ensure accurate selection and order of images in each manual.
4.2.3. QA Pair Refine
We further refine the question-answer pair formulated in the previous steps following the optimization approach outlined by Zhu et al. (2024b). We leverage GPT-4o to identify and extract evidences from contexts that directly supports the answer, while simultaneously recognizing keywords in the answer and checking their alignment with the extracted evidences. We refine the contradictions in the answers by cross-referencing the original answer and the extracted evidence, and discard irrelevant information to ensure the answer is contextually grounded.
@@ 4.3. Data Quality Check
To ensure a high quality and reliability of MRAMG, we further check the consistency and correctness of QA pairs and their alignment with context-related images. We engage annotators in a structured data quality check process. The process is carried out in three stages:
  • In the first stage, a group of annotators review all QA pairs, assessing their correctness and consistency with both the context and image descriptions. They identify any problematic entries for further revision. Moreover, we also integrate GPT-4o for QA Evaluation, to further enhance the quality assurance process.
  • In the second stage, a separate group of annotators correct any identified issues, including those related to image insertions, question formulations, context alignment, image descriptions, and answer correctness. Specifically, expert annotators are further tasked with a meticulous review and correcting over the MRAMG-Arxiv dataset.
  • In the third stage, a team of reviewers conduct a comprehensive recheck to ensure the overall accuracy of the dataset. This step validates that all corrections have been properly implemented and that the datasets meet the required quality standards.

### v2

# MRAMG-Bench: A Comprehensive Benchmark for Advancing Multimodal Retrieval-Augmented Multimodal Generation

@@ Abstract.
Recent advances in Retrieval-Augmented Generation (RAG) have significantly improved response accuracy and relevance by incorporating external knowledge into Large Language Models (LLMs). However, existing RAG methods primarily focus on generating text-only answers, even in Multimodal Retrieval-Augmented Generation (MRAG) scenarios, where multimodal elements are retrieved to assist in generating text answers. To address this, we introduce the Multimodal Retrieval-Augmented Multimodal Generation (MRAMG) task, in which we aim to generate multimodal answers that combine both text and images, fully leveraging the multimodal data within a corpus. Despite growing attention to this challenging task, a notable lack of a comprehensive benchmark persists for effectively evaluating its performance. To bridge this gap, we provide MRAMG-Bench, a meticulously curated, human-annotated benchmark comprising 4,346 documents, 14,190 images, and 4,800 QA pairs, distributed across six distinct datasets and spanning three domains: Web, Academia, and Lifestyle. The datasets incorporate diverse difficulty levels and complex multi-image scenarios, providing a robust foundation for evaluating the MRAMG task. To facilitate rigorous evaluation, MRAMG-Bench incorporates a comprehensive suite of both statistical and LLM-based metrics, enabling a thorough analysis of the performance of generative models in the MRAMG task. Additionally, we propose an efficient and flexible multimodal answer generation framework that can leverage LLMs/MLLMs to generate multimodal responses. Our datasets and complete evaluation results for 11 popular generative models are available at https://github.com/MRAMG-Bench/MRAMG.
Multimodal Retrieval-Augmented Multimodal Generation; Large Language Model; Multimodal Large Language Model; Evaluation
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730288isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrieval
@@ 1. Introduction
With the evolution of generative models, the shift from traditional LLMs to Multimodal Large Language Models (MLLMs) has facilitated a more comprehensive understanding by integrating both textual and visual inputs (Wang et al., 2024b, a).
Correspondingly, Multimodal Retrieval-Augmented Generation (MRAG) (Chen et al., 2022) builds upon traditional RAG by incorporating both textual and visual information, thereby enhancing the quality of generated responses through the image understanding capabilities of MLLMs. However, while the retrieval process incorporates multimodal information, current MRAG methods primarily focus on multimodal input and text-based output. Despite their advanced image comprehension capabilities, MLLMs frequently exhibit hallucinations when converting visual inputs into textual outputs (Bai et al., 2024), which raises a natural question:
Why not present the images directly instead of describing their content?
@@
  • Images as a catalyst for richer answer: “A picture is the icing on the cake.” In applications such as tourist spot descriptions, integrating textual and visual content can significantly enhance the quality of the generated response.
@@
To address this critical demand, we introduce the Multimodal Retrieval-Augmented Multimodal Generation (MRAMG) task, which aims to fully leverage the rich multimodal information within a corpus by generating integrated text-and-image answers (see Section 3 for the formal definition). However, current benchmarks for evaluating this critical task suffer from a notable lack of suitable datasets and scientifically rigorous evaluation metrics. This significant limitation hinders progress in this emerging field.
@@
To bridge this gap, we introduce MRAMG-Bench, a novel benchmark designed to evaluate the MRAMG task comprehensively. MRAMG-Bench consists of six carefully curated English datasets, comprising 4,346 documents, 14,190 images, and 4,800 QA pairs, sourced from three domains—Web, Academia, and Lifestyle—across seven distinct data sources. Furthermore, MRAMG-Bench introduces unique challenges, such as hierarchical difficulty levels and order-based reasoning, which rigorously test the reasoning capabilities of both LLMs and MLLMs. To the best of our knowledge, MRAMG-Bench is the first benchmark that challenges models to autonomously determine the number, selection, and ordering of images in their responses, effectively simulating the complex scenarios encountered in real-world user interactions.
This benchmark is systematically constructed through a multi-stage workflow with both automatic construction of GPT-4o and meticulous formulation of human annotators, as illustrated in Figure 2.
@@
Traditional evaluations of multimodal answers often rely on subjective metrics, such as image-text consistency, which are assessed through LLM-based judgments or human evaluations (Zhang et al., 2023; Zhu et al., 2024a). However, these methods suffer from high subjectivity or evaluator inconsistency. To mitigate these issues, we introduce a statistically grounded evaluation framework that systematically assesses both retrieval and generation performance. By integrating objective statistical metrics with LLM-based assessments, our framework delivers a comprehensive, multi-dimensional evaluation of multimodal answers, ensuring a fair and rigorous assessment process. Additionally, we also propose a general multimodal answer generation framework that integrates rule-based and model-based approaches. This flexible and scalable solution enables both LLMs and MLLMs to generate interleaved text-image responses.
Our main contributions can be summarized as follows:
@@
  • We formulate the MRAMG task and introduce MRAMG-Bench, a benchmark consisting of 4,800 human-annotated instances across multiple domains, featuring hierarchical difficulty levels and order-based reasoning challenges.
  • To ensure rigorous evaluation, we propose a robust, statistically grounded set of metrics that assess multimodal responses across various performance dimensions.
@@
  • We design an efficient and flexible generation framework that leverages LLMs/MLLMs to generate multimodal answers, combining rule-based and model-based approaches.
@@ 3. Task Formulation
In this section, We formally define the task of Multimodal Retrieval-Augmented Multimodal Generation (MRAMG), focusing on constructing multimodal answers $\mathcal{A}$ based on a given text-form user query $q$ and its associated multimodal information $\mathcal{D}$. See Figure 1 for an illustration.
@@ 4. Dataset Construction
Figure 2. The MRAMG-Bench construction pipeline consists of three stages: (1) Data Selection and Preprocessing, where data is collected, cleaned, and preprocessed; (2) QA Generation and Refinement, involving the formulation and refinement of QA pairs; (3) Data Quality Check, where annotators and experts conduct a three-stage review to ensure high-quality benchmarks.
In this section, we present the construction details of MRAMG-Bench. It consists of three domains: Web, Academia and Lifestyle, which respectively contains web-based datasets—MRAMG-Wit, MRAMG-Wiki, MRAMG-Web, and an academic dataset-MRAMG-Arxiv, and two lifestyle datasets-MRAMG-Recipe, MRAMG-Manual. The process of creating datasets in different domains can be divided into three stages: (1) Data Selection and Preprocessing, (2) QA Generation and Refinement, (3) Data Quality Check. The overview of this process is shown in Figure 2.
@@ 4.1. Data Selection and Preprocessing
We undertake three steps during this stage: (1) Data Collection, (2) Data Filtering and Context Enhancement, and (3) Data Refinement.
4.1.1. Data Collection
@@
This category includes Wikipedia pages and articles, derived from Wit dataset (Srinivasan et al., 2021), WikiWeb2M dataset (Burns et al., 2023), and WebQA dataset (Chang and Bisk, 2022). These datasets typically feature a rich integration of text and images. We select val sets and test sets from Wit dataset (Srinivasan et al., 2021), and sample 82k entries from the 100k entries in the test set of WikiWeb2M dataset (Burns et al., 2023) , and choose img-posFacts set in WebQA dataset (Chang and Bisk, 2022), as the initial collection of MRAMG-Wit dataset, MRAMG-Wiki dataset and MRAMG-Web dataset. These datasets primarily consist of single-image documents with relatively short length. Consequently, we categorize the three Web Data datasets as level 1: easy (difficulty level), where models can easily handle our task in this domain due to simple scenarios.
Academic Data.
@@
Datasets in this domain feature in fine-integration of sequences of operation images and instruction texts. These datasets are thoroughly applicable to practical scenarios encountered in users’ daily lives, which are collected from RecipeQA (Yagcioglu et al., 2018), ManualsLib [^1] https://www.manualslib.com and the Technical Dataset [^2] https://www.kaggle.com/datasets/ahmedhgabr/technical-illustration/data. We select the evaluation set from RecipeQA (Yagcioglu et al., 2018), and select 40 categories of instruction manuals with rich text-image content for subsequent construction of MRAMG-Recipe dataset and MRAMG-Manual dataset, respectively. We define Lifestyle Data as level 3: difficult, as they exhibit a significantly higher density of images and encompass diverse problem types, including single-image, no-image, and multi-image answers. Moreover, these datasets pose a significant challenge for models in determining the order of image insertion.
4.1.2. Data Filtering and Context Enhancement
@@
For Academic Data, we engage ten graduate students specializing in computer science, with expertise in fields such as Natural Language Processing (NLP) and Computer Vision (CV), to serve as expert annotators. Each annotator is assigned papers within their specific domain of expertise, ensuring precise and contextually relevant annotations. The annotators meticulously read the selected papers, identify and extract textual content and corresponding figures that are most suitable for creating interleaved text-image contexts. This ensures accurate alignment of text and images in contexts.
For Lifestyle Data, in MRAMG-Recipe dataset, each recipe contains multiple step texts and each step text contains multiple images. We first combine multiple images into one image for each recipe step text and insert the combined image at the end of each step text. We then aggregate all the recipe step texts as an entire image-text interleaved recipe context. In MRAMG-Manual dataset, we employ MinerU (Wang et al., 2024c) to parse the PDF manuals into markdown format, preserving the original contextual order of images and texts within each manual. A team of annotators is tasked with removing sensitive or irrelevant information, adjusting image placements within the content, and correcting any errors in the text.
4.1.3. Data Refinement
@@ 4.2. QA Generation and Refinement
This stage can be further divided into three steps: (1) Question Generation, (2) Answer Generation, (3) QA Pair Refinement.
@@ 4.2.1. Question Generation
For MRAMG-Web dataset, we retain the questions contained in original dataset. As for MRAMG-Wiki, MRAMG-Wit and MRAMG-Recipe datasets, we leverage GPT-4o to generate high-quality, context-specific questions tailored to the provided text and images. Moreover, the questions in MRAMG-Manual are generated manually by human annotators. As for MRAMG-Arxiv dataset, we engage expert annotators to meticulously formulate questions. Overall, we adhere to the following two criteria: (1) The questions are constructed based on the corresponding contexts. (2) The questions should be natural and practical, with the potential to be effectively answered through the integration of text and images.
@@ 4.2.2. Answer Generation
For Web Data, given the generated or original questions $Q$ and corresponding context $C$, we generate image-text interleaved answers $A$ using GPT-4o following a CoT reasoning strategy (Wei et al., 2022): (1) Question Validity Assessment, we first ensure that valid questions $ValQ$ can be directly answered by contexts $C$ and remove invalid questions $InvalQ$. (2) Evidence Extraction, we then extract evidence $E$ from contexts $C$ to substantiate the answer $A$ for later answer construction. (3) Answer Construction, we construct a highly reliable, accurate and coherent answer based on valid questions $ValQ$ and evidence $E$. (4) Image Integration we finally enrich the constructed answers $A$ with a series of images integrated in the optimal positions. For Academic Data, expert annotators review existing questions, create answers incorporating images or tables, and integrate visual elements to ensure clear, multimodal comprehension of the content. For Lifestyle Data, in MRAMG-Recipe dataset, there are two types of questions: questions related to the entire recipe (recipe-specific) and questions only related to a specific step (step-specific). We combine all the step texts as the answers of recipe-specific questions and use answers generated by GPT-4o as the answers of step-specific questions. We manually insert images at the end of answer texts. In MRAMG-Manual dataset, answers are meticulously formulated by annotators, to ensure accurate selection and order of images in each manual.
4.2.3. QA Pair Refinement
We further refine the QA pairs formulated in the previous steps following the optimization approach outlined by Zhu et al. (2024b). We leverage GPT-4o to extract supporting evidence from contexts and verify its alignment with keywords in the answer. We resolve the contradictions in the answer by cross-referencing the original answer and the extracted evidence, and discard irrelevant information to ensure the answer is contextually grounded.
@@ 4.3. Data Quality Check
To ensure a high quality and reliability of MRAMG, we further check the consistency and correctness of QA pairs and their alignment with context-related images. We engage annotators in a structured data quality check process. The process is carried out in three steps:
  • In the first step, a group of annotators review all QA pairs, assessing their correctness and consistency with both the context and image descriptions. They identify any problematic entries for further revision. Moreover, we also integrate GPT-4o for QA Evaluation, to further enhance the quality assurance process.
  • In the second step, a separate group of annotators correct any identified issues, including those related to image insertions, question formulations, context alignment, image descriptions, and answer correctness. Specifically, expert annotators are further tasked with a meticulous review and correcting over the MRAMG-Arxiv dataset.
  • In the third step, a team of reviewers conduct a comprehensive recheck to ensure the overall accuracy of the dataset. This step validates that all corrections have been properly implemented and that the datasets meet the required quality standards.